Four years ago Katie was twice as old as Anne was then. in 6 years, Anne will be the same age that katie is now. How old is each now ?

I have two equattions of Katie= X, y= Anne
2X-4= y
x+6 = 2y - 4

Answer by drj(1380)

Four years ago Katie was twice as old as Anne was then. In 6 years, Anne will be the same age that Katie is now. How old is each now ?

I have two equations of Katie= X, y= Anne

2X-4= y
x+6 = 2y - 4

You're in the right track and good try in getting two linear equations....I'll check your work by the following steps.

Step 1. Let x be Katie's age and y be Anne's age.

Step 2. Let x-4 be Katie's age 4 years ago and y-4 be Anne's age four years ago.

Step 3. Then, x-4= 2(y-4) since four years ago Katie was twice as old as Anne was then. Simplifying yields x-2y=-4

Step 4. Let y+6 be the age of Anne.

Step 5. Then y+6=x since in 6 years, Anne will be the same age that Katie is now. Simplifying yields x-y=6

Step 6. The system of equations from Steps 3 and 5 are

x-2y=-4 Equation A
x-y-=6 Equation B

OR multiplying Equation B by -1 yields

x-2y=-4 Equation A
-x+y=-6 Equation B1

Adding yields

x-x-2y+y=-4-6

or y=10 and x=10+6=16

Substituting x=16 and y=10, into Steps 3 and 5 yields true statements.

Step 7. ANSWER: Kate is 16 years old and Anne is 10 years old.

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J